OpenStudy (bowtie_man123):

*WILL GIVE MEDDLE* Set up the following word problem as a proportion and solve. Then write your answer as a complete sentence. A health department estimates that 8 vials of malaria serum will treat 100 people. At this rate, how many vials are required to treat 175 people?

OpenStudy (bowtie_man123):

@Jack_Prism @Kikuo

OpenStudy (kikuo):

8/100=175/x

OpenStudy (kikuo):

100x175=8x

OpenStudy (flower-girl):

Hello 8 vials = 100 people x vials = 175 people 175*8=100x x=175*8/100 When you do that what do you get?

OpenStudy (kikuo):

Never mind. Flower-girl is correct. I setup my proportion incorrectly. The answer is 14.

OpenStudy (flower-girl):

:)

OpenStudy (bowtie_man123):

i sorry i just did it and it is 14 now i jut have convert it into a sentence

OpenStudy (flower-girl):

oh okay...

OpenStudy (bowtie_man123):

i am sorry opps lol i typed i sorry

OpenStudy (bowtie_man123):

could you guys or girls or 1 guy one girl help me with the sentance

OpenStudy (flower-girl):

What kind of sentence...

OpenStudy (bowtie_man123):

nevermind i got it but thanks and i gave you your meddle and fanned you

OpenStudy (flower-girl):

Thanks! :)

OpenStudy (bowtie_man123):

np